# Dam breaks in the discrete nonlinear Schrödinger equation

### v1

@@ III Linear dispersion relation of two-phase waves
±$\pm 4\sqrt{\beta}\sin\left(\frac{kh}{2}\right)\sqrt{\cos(\kappa h)}\sqrt{\beta\sin^{2}\left(\frac{kh}{2}\right)\cos(\kappa h)+\frac{A^{2}}{2}},$
@@ V Dynamics of weak DSW
Figure 3: Snapshots of small amplitude DNLS DSW emitted at $t=1000$ for $u_{+}=0.9$, as $\beta$ varies from $0.1$ to $0.6$. Here, in (a) $\beta=0.1$, and we obtain DSW in a regime with negative dispersion curvature $\tilde{\alpha}_{3}<0$. Thereafter, as $\beta$ is increased, in steps of $0.1$, we observe traveling DSW in (B) and (C) which correspond to $\beta=0.2$, $0.3$ respectively and crossover DSW for (D), corresponding to $\beta=0.4$ and finally DSW in a regime with positive dispersion curvature in (E), (F) [$\beta=0.5$, $0.6$ respectively], whose waveforms are significantly influenced by higher-order dispersive effects.
@@ VI.2 Wave patterns in the weak coupling limit
In this section, we unravel the characteristics of wave patterns that emerge from a family of dam break problems with $u_{-}=1$, and a representative $\beta=0.1$ while $u_{+}$ varies in the interval $[-1,1)$. I.e., complementarily to the considerations in the previous subsection (where the jump height was maintained fixed while the coupling $\beta$ was varied), here we maintain a fixed coupling, while changing the jump height. In particular, we attempt to study the emitted DSW patterns within the framework of simple wave DSW theory, and we utilize the failure of the theory to corroborate the numerical observations.
@@
To obtain some analytical insight into the bright excitation (observed frequently for $u_{+}<0.5$), we isolate a representative small amplitude ($|a|\ll u_{+}$) mode emitted from a Riemann problem characterized by $u_{+}=0.45$ and $\beta=0.1$. The pulsating small amplitude excitation is reminiscent of oscillatory Kawahara solitons [61, 53]. We perform such a fit of a numerically computed Kawahara soliton [62] in the co-traveling frame $\tilde{\xi}=x-ct$ (Fig. 14). Here, two singificantly different realizations of the traveling breather are compared to the Kawahara soliton (c.f. Eq. (35) whose coefficients defined by the background $u_{+}$). We find that the Kawahara soliton approximates the traveling “breather” reasonably well with the relative error across snapshots limited to $4\%$. Several such small-amplitude traveling “bright” breathers appear to be well approximated by the Kawahara solitons. The larger-amplitude (and faster) variants of such bright excitations have a distinct box-type profile (cf. Fig. 6 (D)). To reveal some of the properties of this box-state, we follow its center of mass using the formula $x_{\rm COM}$ $=$ $\bigg{(}\sum_{k=1}^{N}\frac{1}{\sqrt{\beta}}k|u_{k}|^{2}\bigg{)}$ $\bigg{(}\sum_{k=1}^{N}|u_{k}|^{2}\bigg{)}^{-1}$. A linear fit to the center of mass function unveils its approximate speed (c.f. Fig. 15 (A)) to be $\approx 1.047$. The periodic pulsations of this breather in the co-moving frame can be probed by computing the Fast Fourier transform of the periodic time series $x_{\rm COM}-v_{\rm COM}t$. This reveals the dominant frequency $\approx 1/3$ (Fig. 15 (C)). This calculation is verified directly by time evolution of the waveform up to $t=3$, which reveals the waveform reverts to its original shape (Fig. 15 (D)). The complete characterization of this family of breathers, most notably in the context of the observed box waves is yet another intriguing feature worthy of consideration for further work. Indeed, this is technically a rather intriguing topic as it involves the potential for traveling states within the DNLS model, a feature that has been long debated in numerous earlier works [63, 64].
@@
We have now set the stage for the examination of the corresponding dam breaks in the weak-coupling limit. Snapshots of the dynamics at $t=2000$ have been shown in Fig. 17. For $u_{+}=-1$ (Fig. 17 (A)), besides the intersite DS formed at the central sites, counterpropagating “bright” moving (indeed, practically traveling) breathers are generated. The observed moving structures here are nearly solitary wave-like, and in the given dispersion regime $\tilde{\alpha}_{3}(|u_{+}|,0)<0$ may be approximated by KdV (Eq. (29)) solitons of the form $|u_{+}|+a_{s}{\rm sech}^{2}\left(\sqrt{\frac{a_{s}}{12}}\sqrt{-\frac{\tilde{\beta}}{\tilde{\alpha_{3}}}}\epsilon^{1/2}(x-(\tilde{\alpha}_{1}+V_{s}\tilde{\beta}\epsilon)t)\right)$. We find that these KdV solitons for a given amplitude are narrower in extent. Perhaps the inclusion of higher-order dispersion could lead to a closer agreement. We have also seen that the dynamically observed DS compares well with a solution obtained by solving a boundary value problem for the discrete NLS utilizing Newton iterations [65]. Excellent initial guesses for the Newton iterations leading to a stationary DS can be generated in the weak-coupling limit by considering the perturbation expansions $v_{n}=v^{(0)}_{n}+\beta v^{(1)}_{n}+\beta^{2}v^{(2)}_{n}+\cdots$ with the zeroth-order solution arising in the anticontinuum limit (cf. [66, 65]).
@@ VI.3.2 Kinks
Next, we turn to the leading order corrections to the stability spectrum near the anti-continuum limit. In the anti-continuum limit ($\beta=0$), the eigenvalues of the stability matrix (Eq. (40)) can be obtained (explicitly) either by looking at the zero and nonzero sites, which are associated with $\lambda=\pm i\mu$ and $\lambda=0$, respectively. Now, in the finite (but small) $\beta$ case, an approximation to the spectrum can be obtained. At first, we note the emergence of the continuous spectrum given by $-i\sqrt{16\beta^{2}+8\beta u_{-}^{2}}\leq\lambda\leq i\sqrt{16\beta^{2}+8\beta u_{-}^{2}}$. Furthermore, we consider the perturbations of the eigenvalue ±$\pm i\mu$ as we depart gradually from the anti-continuum limit. For this, we consider the $4\times 4$ diagonal block associated with the sites $n=-1,0$.
@@
For completeness, we also derived the higher-order correction to the spectrum of the stability operator in (40) by analyzing contributions from the sites $n=-2,-1,0,1$ , retaining terms up to $\mathcal{O}(\beta^{2})$. This refinement yields a slightly more accurate prediction for the onset of the Hamiltonian-Hopf bifurcation, at $\tilde{\beta}_{c}\approx 0.0684\mu$. Besides, these calculation predict the presence of four embedded eigenvalues within the continuous spectrum when $\beta<\tilde{\beta}_{c}$. I.e., presumably in extending the calculation to more sites, in addition to the eigenvalue pair of interest, we are progressively accounting for modes within the continuous spectrum.
@@
An extensive stability analysis suggests a dependence of the stability spectra on the number of sites present in the lattice, for $\beta>\tilde{\beta_{c}}$ (analogously to the case of dark solitons [68]). However, for $\beta$ close to the threshold $\beta_{c}$ the instability growth rates are minimal. For example, with $1200$ sites, the kink characterized by $\beta=0.1$ (and $\mu=1$), has ${\rm max}({\mathop{\mathrm{Re}}(\lambda)})\approx 1.22\times 10^{-4}$ (c.f. Fig. 22, where the growth rate is verified through direct numerical simulations). Thus, the kinks are observed resulting from the Riemann problems for reasonably long times ($t=1000$) in Fig. 6 (F). However, for Riemann problems with increasing $\beta$ these structures are subject to progressively more pronounced instability growth rates. This can be leveraged to explain why they are not observed to persist even for short times therein (c.f. Fig. 23).
@@
Figure 21: (A) Kink computed numerically (in red) and found analytically in the small $\beta$ limit (in blue), for $\beta=1/20$, (B) The numerically obtained spectrum (in red) compared to the analytical formulas for the same $\beta=1/20$ (discrete eigenvalues: blue stars, continuous spectrum: blue dots), (C) Kink computed numerically (in red) and found analytically in the small $\beta$ limit (in blue), for $\beta=1/14$, which lies in the regime just past the Hopf bifurcation point derived analytically to $\approx 0.0771$, (D) The eigenspectrum (analytical: blue versus numerical: red) shown for the same $\beta=1/14$.
Figure 22: The instability growth rate observed for the kink characterized by $\mu=1$ and $\beta=0.1$, and perturbed with the unstable eigenmode whose eigenvalue ${\rm Re}(\lambda)=1.22\times 10^{-4}$. A logarithm-fit based on the growth rate prediction from linear stability analysis shows good agreement with direct numerical simulation.

### v2

@@ III Linear dispersion relation of two-phase waves
$\pm 4\sqrt{\beta}\sin\left(\frac{kh}{2}\right)\sqrt{\cos(\kappa h)}\sqrt{\beta\sin^{2}\left(\frac{kh}{2}\right)\cos(\kappa h)+\frac{A^{2}}{2}},$
@@ V Dynamics of weak DSW
Figure 3: Snapshots of small amplitude DNLS DSW emitted at $t=1000$ for $u_{+}=0.9$, as $\beta$ varies from $0.1$ to $0.6$. Here, in (a) $\beta=0.1$, and we obtain DSW in a regime with negative dispersion curvature $\tilde{\alpha}_{3}<0$. Thereafter, as $\beta$ is increased, in steps of $0.1$, we observe traveling DSW in (B) and (C) which correspond to $\beta=0.2$, $0.3$ respectively and crossover DSW for (D), corresponding to $\beta=0.4$ and finally DSW in a regime with positive dispersion curvature in (E), (F) $\beta=0.5$, $0.6$ respectively], whose waveforms are significantly influenced by higher-order dispersive effects.
@@ VI.2 Wave patterns in the weak coupling limit
In this section, we unravel the characteristics of wave patterns that emerge from a family of dam break problems with $u_{-}=1$, and a representative $\beta=0.1$ while $u_{+}$ varies in the interval [1,1)[-1,1). I.e., complementarily to the considerations in the previous subsection (where the jump height was maintained fixed while the coupling $\beta$ was varied), here we maintain a fixed coupling, while changing the jump height. In particular, we attempt to study the emitted DSW patterns within the framework of simple wave DSW theory, and we utilize the failure of the theory to corroborate the numerical observations.
@@
To obtain some analytical insight into the bright excitation (observed frequently for $u_{+}<0.5$), we isolate a representative small amplitude ($|a|\ll u_{+}$) mode emitted from a Riemann problem characterized by $u_{+}=0.45$ and $\beta=0.1$. The pulsating small amplitude excitation is reminiscent of oscillatory Kawahara solitons [61, 53]. We perform such a fit of a numerically computed Kawahara soliton [62] in the co-traveling frame $\tilde{\xi}=x-ct$ (Fig. 14). Here, two singificantly different realizations of the traveling breather are compared to the Kawahara soliton (c.f. Eq. (35) whose coefficients defined by the background $u_{+}$). We find that the Kawahara soliton approximates the traveling “breather” reasonably well with the relative error across snapshots limited to $4\%$. Several such small-amplitude traveling “bright” breathers appear to be well approximated by the Kawahara solitons. The larger-amplitude (and faster) variants of such bright excitations have a distinct box-type profile (cf. Fig. 6 (D)). To reveal some of the properties of this box-state, we follow its center of mass using the formula $x_{\rm COM}$ == $\bigg(\sum_{k=1}^{N}\frac{1}{\sqrt{\beta}}k|u_{k}|^{2}\bigg)$ $\bigg(\sum_{k=1}^{N}|u_{k}|^{2}\bigg)^{-1}$. A linear fit to the center of mass function unveils its approximate speed (c.f. Fig. 15 (A)) to be $\approx 1.047$. The periodic pulsations of this breather in the co-moving frame can be probed by computing the Fast Fourier transform of the periodic time series $x_{\rm COM}-v_{\rm COM}t$. This reveals the dominant frequency $\approx 1/3$ (Fig. 15 (C)). This calculation is verified directly by time evolution of the waveform up to $t=3$, which reveals the waveform reverts to its original shape (Fig. 15 (D)). The complete characterization of this family of breathers, most notably in the context of the observed box waves is yet another intriguing feature worthy of consideration for further work. Indeed, this is technically a rather intriguing topic as it involves the potential for traveling states within the DNLS model, a feature that has been long debated in numerous earlier works [63, 64].
@@
We have now set the stage for the examination of the corresponding dam breaks in the weak-coupling limit. Snapshots of the dynamics at $t=2000$ have been shown in Fig. 17. For $u_{+}=-1$ (Fig. 17 (A)), besides the intersite DS formed at the central sites, counterpropagating “bright” moving (indeed, practically traveling) breathers are generated. The observed moving structures here are nearly solitary wave-like, and in the given dispersion regime $\tilde{\alpha}_{3}(|u_{+}|,0)<0$ may be approximated by KdV (Eq. (29)) solitons of the form $|u_{+}|+a_{s}{\rm sech}^{2}\left(\sqrt{\frac{a_{s}}{12}}\sqrt{-\frac{\tilde{\beta}}{\tilde{\alpha_{3}}}}\epsilon^{1/2}(x-(\tilde{\alpha}_{1}+V_{s}\tilde{\beta}\epsilon)t)\right)$. We find that these KdV solitons for a given amplitude are narrower in extent. Perhaps the inclusion of higher-order dispersion could lead to a closer agreement. We have also seen that the dynamically observed DS compares well with a solution obtained by solving a boundary value problem for the discrete NLS utilizing Newton iterations [65]. Excellent initial guesses for the Newton iterations leading to a stationary DS can be generated in the weak-coupling limit by considering the perturbation expansions $v_{n}=v^{(0)}_{n}+\beta v^{(1)}_{n}+\beta^{2}v^{(2)}_{n}$$\cdots$ with the zeroth-order solution arising in the anticontinuum limit (cf. [66, 65]).
@@ VI.3.2 Kinks
Next, we turn to the leading order corrections to the stability spectrum near the anti-continuum limit. In the anti-continuum limit ($\beta=0$), the eigenvalues of the stability matrix (Eq. (40)) can be obtained (explicitly) either by looking at the zero and nonzero sites, which are associated with $\lambda=\pm i\mu$ and $\lambda=0$, respectively. Now, in the finite (but small) $\beta$ case, an approximation to the spectrum can be obtained. At first, we note the emergence of the continuous spectrum given by $i\sqrt{16\beta^{2}+8\beta u_{-}^{2}}\leq\lambda\leq i\sqrt{16\beta^{2}+8\beta u_{-}^{2}}$. Furthermore, we consider the perturbations of the eigenvalue $\pm i\mu$ as we depart gradually from the anti-continuum limit. For this, we consider the $4\times 4$ diagonal block associated with the sites n=1,0n=-1,0.
@@
For completeness, we also derived the higher-order correction to the spectrum of the stability operator in (40) by analyzing contributions from the sites n=2,1,0,1n=-2,-1,0,1 , retaining terms up to $\mathcal{O}(\beta^{2})$. This refinement yields a slightly more accurate prediction for the onset of the Hamiltonian-Hopf bifurcation, at $\tilde{\beta}_{c}\approx 0.0684\mu$. Besides, these calculation predict the presence of four embedded eigenvalues within the continuous spectrum when $\beta<\tilde{\beta}_{c}$. i.e., presumably in extending the calculation to more sites, in addition to the eigenvalue pair of interest, we are progressively accounting for modes within the continuous spectrum. A sharp estimate of the (in)stability threshold can be obtained by tracking the collision of the continuous spectral bands associated with the zero background ($(1-4\beta)\leq|\lambda|\leq 1$) and the non-zero background ($|\lambda|\leq\sqrt{16\beta^{2}+8\beta u_{-}^{2}}$) of the kink, which carry opposite Krein signatures. For $u_{-}=1$, this is seen to occur for $\tilde{\beta}_{c}=1/16=0.0625$, determined by $\sqrt{16\beta^{2}+8\beta}=1-4\beta$. Checking this against the numerically predicted threshold (i.e. examining the parametric ($\beta$) variation of eigenvalues of Eq. (40)), we find that for $N=800$ this transition to complex quartets with associated real part $\sim 10^{-6}$ occurs at $\beta\approx 0.0629$.
@@
An extensive stability analysis suggests a dependence of the stability spectra on the number of sites present in the lattice, for $\beta>\tilde{\beta_{c}}$ (analogously to the case of dark solitons [68]). However, for $\beta$ close to the threshold $\beta_{c}$ the instability growth rates are minimal. For example, with $1200$ sites, the kink characterized by $\beta=0.1$ (and $\mu=1$), has ${\rm max}({\mathop{\mathrm{Re}}(\lambda)})\approx 1.22\times 10^{-4}$. Thus, the kinks are observed resulting from the Riemann problems for reasonably long times ($t=1000$) in Fig. 6 (F). However, for Riemann problems with increasing $\beta$ these structures are subject to progressively more pronounced instability growth rates. This can be leveraged to explain why they are not observed to persist even for short times therein (c.f. Fig. 22).
To verify the linear stability predictions associated with the unstable kinks, we seeded their profiles with the maximally unstable eigenmode (i.e. $u_{n}=v_{n}+\epsilon(p_{n}+q_{n}^{*})$, $\epsilon\ll 1$) and tracked their complex-valued growth rate numerically by monitoring ${\rm max}_{n}||u_{n}(t)|-v_{n}|$. To corroborate this against the analytical result, we perform a numerical fit (for $\epsilon$) in the dynamical linear stability prediction $u_{n}=\exp(-i\mu t)[v_{n}+\epsilon(p_{n}e^{\lambda t}+q_{n}^{*}e^{\lambda^{*}t})]$ where $\lambda$ is the eigenvalue associated with the maximally unstable eigenmode. Note that not only do we capture the exponential growth (in time) ${\rm Re}(\lambda)\approx 0.0302$, but also capture the oscillatory part of the instability ${\rm Im}(\lambda)\approx 0.258$ accurately (see also Fig. 23 for the case $\beta=0.18$).
@@
Figure 21: (A) Kink computed numerically (in red) and found analytically in the small $\beta$ limit (in blue), for $\beta=1/20$, (B) The numerically obtained spectrum (in red) compared to the analytical formulas [perturbation theory result] for the same $\beta=1/20$ (discrete eigenvalues: blue stars, continuous spectrum: blue dots), (C) Kink computed numerically (in red) and found analytically in the small $\beta$ limit (in blue), for $\beta=1/14$, which lies in the regime across the Hopf bifurcation point, (D) The eigenspectrum (analytical: blue versus numerical: red) shown for the same $\beta=1/14$.
@@ VI.3.3 Vacuum dam breaks for intermediate $\beta$
Figure 23: The growth of the (oscillatory) instability mode associated with a kink characterized by $\beta=0.18$ and $N=800$ (number of sites). The extracted growth rate (numerically, see red solid line) agrees with the analytical prediction $\underset{\lambda_{i}}{\rm max}[{\rm Re}(\lambda)]\approx 0.0302$ to within $0.2\%$. Likewise, the imaginary part of the same eigenvalue ($\approx 0.258$) accurately captures the oscillatory nature of the unstable mode, exhibiting the same level of relative accuracy, as evidenced by the agreement between the (dynamical) quasi-analytical (blue dots) and numerical (black solid) curves in the time interval $t\in[100,600)$.